OpenStudy (anonymous):

help please! (: The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected!

hero (hero):

If x = 1st integer x + 1 = 2nd integer x + 2 = 3rd integer Then: x + (x + 1) + (x + 2) = 267 x + x + 1 + x + 2 = 267 x + x + x + 1 + 2 = 267 3x + 3 = 267 Solve for each integer.

OpenStudy (anonymous):

is it x=90?

hero (hero):

Not exactly

OpenStudy (anonymous):

i don't get it :/

OpenStudy (anonymous):

According to Hero...but 90 'are' not 3 consecutive integers.

hero (hero):

@sylinan, why don't you show your steps for how you got 90

OpenStudy (anonymous):

X+X-1+X-2 =267 3X -3 = 267 3X = 270 X = 270/3 X = 90

hero (hero):

Okay, I see

hero (hero):

But, you should do it the way I have shown next time.

hero (hero):

So what are the other two integers?

hero (hero):

@Tommy, you're way off bro.

OpenStudy (anonymous):

im so lost);

OpenStudy (anonymous):

5*6*7=210 6*7*8=336 QED no integers solve this question.

hero (hero):

@sylinan, what are the other two integers?

OpenStudy (anonymous):

90 x 91 x 92.......

hero (hero):

@Tommy, those are not the integers

OpenStudy (anonymous):

im not sure how to solve it

hero (hero):

@sylinan, you already found x = 90

OpenStudy (anonymous):

Oh dear I'm getting confused with the sum & the product.... sorry guys.

OpenStudy (anonymous):

but i thought it was wrong

hero (hero):

I just said you did it differently @sylinan. Afterwards, I asked you to continue finding the other two integers.

OpenStudy (anonymous):

oh.. my bad

OpenStudy (anonymous):

I'm wrong.. you know 'x' now find the other 2 integers the other two 'x's.

hero (hero):

Nobody is going to give you the answer. Especially not @T0mmy

OpenStudy (anonymous):

Now even I know the answer Hero... merely a xmas mistake...cheers. ;D

OpenStudy (anonymous):

haha i know okay x + x + 1 + x +2 = 267 3x + 3 = 267 3x = 267 - 3 3x = 264 ?

hero (hero):

You can do it either way @sylinan

hero (hero):

Both formulas work. All you need to do is figure out all three integers.

hero (hero):

You found out that 90 is one of them. What are the other two?

OpenStudy (anonymous):

I did lots og Guinness theen 7s angd got to the answert

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

Thanks T0m goood nioght.